(1 point) A die s continuously rolled until the total sum of all rolls exceeds 175 What is the probability that at least 55 rolls are necessary? 0.1423

Please show work so i can understand it. Thank you!
Answer #1

for a single roll each outcomes has equal probbaility =1/6

x f(x) xP(x) x2P(x)
1 1/6 0.167 0.167
2 1/6 0.333 0.667
3 1/6 0.500 1.500
4 1/6 0.667 2.667
5 1/6 0.833 4.167
6 1/6 1.000 6.000
total 3.500 15.167
E(x) =μ= ΣxP(x) = 3.5000
E(x2) = Σx2P(x) = 15.1667
Var(x)=σ2 = E(x2)-(E(x))2= 2.9167
std deviation=         σ= √σ2 = 1.7078

hence for 54 rolls ; expected total =54*3.5=189

and std deviation =1.7078*sqrt(54)=12.55

therefore P(that for 55 rolls sum is at most 175)= P(X<175)=P(Z<(175.5-189)/12.55)=P(Z<-1.08)=0.1401

( please revert for any clarifcation)